OpenStudy (anonymous):

Bob and Mary is looking at Johns paper and see matrix multiplication problem He know that that Johns paper is [0,-1][2,4,6] [1,0] [1,6,0]

OpenStudy (anonymous):

A. rotating a triangle 90º clockwise B. rotating a triangle 90º counterclockwise C. reflecting a triangle in the line y = x D. reflecting a triangle in the x-axis

OpenStudy (amistre64):

run each column from the left thru the right matrix to define new positions

OpenStudy (amistre64):

i cant tell right form left today :/

OpenStudy (amistre64):

take 2,1 2*0+1*-1 = 0 - 1 = -1 1 0 = 2+ 0 = 2 2,1 -> -1,2

OpenStudy (amistre64):